Challenger Factory produces two similar products - regular widgets and deluxe widgets. The total plant overhead budget is $625,400 with 369,500 estimated direct labor hours. It is further estimated that deluxe widget production will need 3 direct labor hours for each unit and regular widget production will require 5 direct labor hours for each unit.

Using the single plantwide factory overhead rate with an allocation base of direct labor hours, how much factory overhead will Challenger Factory allocate to regular widget production if budgeted production for the period is 75,000 units and actual production for the period is 103,900 units?
Answer #1

Solution:

Plantwide overhead rate = Estimated overhead cost/ estimated direct labor hours = $625400/369500 = $1.692558 per DLH

Overhead allocated to Regular widgets = direct labor hours for regular widgets * plantwide overhead rate

= (103900*5) *$1.692558 = $879,284 (rounded)